# Catalog of outbursts of neutron star LMXBs

### v1

@@ Abstract
Many X-ray binaries are transiently accreting. Having statistics on their recurrence times is helpful to address questions related to binary evolution and populations, as well as the physics of binary systems. We compile a catalog of known outbursts of neutron star (identified through bursts or pulsations) low-mass X-ray binaries, until late 2023. Most outbursts are taken from the literature, but we also identify some outbursts from public X-ray monitoring lightcurves. We find 109 outbursts not previously identified in the literature; most are from the frequent transients GRS 1747-312 and the Rapid Burster MXB 1730-335, though we suspect that two outbursts from Liller 1 may be from another transient, besides the Rapid Burster. We also find new outbursts for 10 other systems, and verify substantial quiescent intervals for XMM J174457-2850.3, XMMU J174716.1-281048, and AX J1754.2-2754. Outburst detection has been relatively efficient since 1996 for outbursts above $F_{X}$(2-10)$=3\times 10^{-10}$ ergs/s/cm2. While several systems have many known outbursts, 40 of the 85 systems we track have zero or one recorded outburst between 1996 and 2023. This suggests that faint Galactic Center X-ray binaries may be neutron star X-ray binaries, though we cannot completely rule out the proposition that most neutron star X-ray binaries undergo frequent outbursts below all-sky monitor detection limits.
@@ 1 Introduction
The total number of mass-transferring neutron star low-mass X-ray binaries in the Galaxy is thought to be of the order of $10^{4}$ (Portegies Zwart et al., 1997; Kiel & Hurley, 2006; Jonker et al., 2011; van Haaften et al., 2015). However, only a small fraction of these systems have been detected, typically during outbursts in which their X-ray luminosity increases by a factor of $10^{3}$$10^{7}$ compared to quiescence. Understanding the full population of these systems, and their binary evolution, is helpful for us to understand the physics of binary systems (e.g. Podsiadlowski et al., 2002; Postnov & Yungelson, 2014; Van et al., 2019), and also to understand the products of their evolution, such as millisecond pulsars (Alpar & Shaham, 1985; Tauris & Savonije, 1999; Tauris et al., 2012), which may produce the excess gamma-rays seen from the Galactic bulge (Wang et al., 2005; Gonthier et al., 2018). One question of current interest is the nature of X-ray sources of intermediate luminosity ($L_{X}$(2-10 keV)$\sim 10^{31}-10^{33}$ erg/s) detected near the Galactic Centre (Muno et al., 2005; Hailey et al., 2018; Mori et al., 2021; Maccarone et al., 2022). The objects in this region that show no outbursts have been argued to be mostly black hole LMXBs, based on an argument that neutron star LMXBs do not have long enough recurrence times. Having detailed statistics on X-ray outbursts from neutron stars can be helpful in beginning to answer questions such as these.
@@
A number of catalogs of X-ray binaries have been produced over the years. The catalogs of Liu et al. (2001, 2007) have been particularly well-cited, but are now 17 years old. Ritter & Kolb (2003) have an extensive and updated catalog of X-ray binaries and cataclysmic variables with known periods, but do not include objects without known (or suggested) orbital periods. More recent catalogs have been published of black hole X-ray binaries, one focused on compiling detailed lightcurves (Tetarenko et al., 2016b), and one focusing on multiwavelength studies (Corral-Santana et al., 2016). We are interested in the particular problem of understanding the frequency of X-ray outbursts from neutron star LMXBs, but found that we lacked a catalog to answer this question. The catalog of Avakyan et al. 2023 was released when this work was largely complete, and while it is very useful, it also does not list outbursts. Yan & Yu (2015) and Lin et al. (2019) construct tables of outburst frequency for a number of X-ray binaries with known orbital periods, but these works only include 14 and 19 NS LMXBs respectively. Yan & Yu (2015) do include a detailed listing of the outbursts they include, which is helpful. We identify several more outbursts from 4U 1608-52, XTE J1709-267, MXB 1730-33, XTE J1739-285, 1A 1744–361, SAX J1748.9-2021, and Aquila X-1 in the time range these authors cover, typically fainter outbursts.
@@ 2 Catalog
We omit 42 NS LMXBs in the MINBAR catalog that appear to be persistent. We list these persistent, verified bursting NS XRBs in Table D, in Appendix D.
@@
We omit a few transient bursters; the position of XB 1940-04 (Murakami et al., 1983) is very poorly known. The eccentric X-ray binary Cir X-1 (or 3A 1516-569) has outbursts that reflect its orbital period, not disk outbursts (Murdin et al., 1980). We exclude 1A 1742-289 (or A1742-289), as it is not clear that the burst activity seen from MXB 1743-29 in the Galactic Center in 1976 (Lewin et al., 1976a) can be attributed to this radio-bright (Davies et al., 1976) source (we note that this is definitely not the same object as AX J1745.6-2901; Kennea & Skinner 1996). This leaves 70 sources from the MINBAR catalog.
@@
We add to the MINBAR list the known recent bursters 4U 1730-22, MAXI J1816-195, MAXI J0556-332, MAXI J0911-655, Terzan 6 X-2 (see Appendix § B), and SRGA J144459.2-604207 111as updated by Jean in ’t Zand; https://personal.sron.nl/~jeanz/bursterlist.html.  [^2] We include SRGA J144459.2-604207 in our list for completeness, but otherwise our summary of outbursts stops in 2023. We also add known transient accreting millisecond pulsars that are not known bursters, from the Patruno & Watts (2021) catalog and more recent literature, XTE J0929-314, NGC 6440 X-2, XTE J1751-305, Swift J1756.9-2508, XTE J1807-294, IGR J17494-3030, IGR J16597-3704, and MAXI J1957+032. Finally, we include the transient slow (2.14 Hz) accreting pulsar GRO J1744-28. (We also list the few persistent LMXBs that are pulsars in Table D.) This gives us a total of 85 transient neutron star low-mass X-ray binaries.
@@
We performed literature searches for this list of transients, and identify the known X-ray outbursts from each system. For each outburst, we attempt to determine the peak observed 2-10 keV X-ray luminosity. This will have some significant uncertainty, since not every outburst is caught at its peak, some outbursts are observed only in other energy ranges, and many distances are not well-known. We attempt to use the best-fit spectrum to convert measurements to 2-10 keV fluxes, but we only attempt this with simple power-law spectra (we do not try to convert complex Comptonized models). Where information is lacking, we use the best literature $N_{H}$ estimate and an assumed photon index of 2.1 to make conversions, or use literature estimates of the 2-10 keV flux in mCrab (assuming 1 Crab = $2\times 10^{-8}$ ergs cm-2 s-1). A few bursters have recorded bursts without any recorded outbursts (Gotthelf & Kulkarni, 1997; Cornelisse et al., 2002b); we list these in Table C in Appendix C.
@@
We present the catalog of known outbursts of neutron star LMXBs in Table 1, for which we provide a sample here. The full table is in the electronic version of the journal (or Appendix A for the arXiv version). We include a line for each recorded outburst, with a literature reference for each outburst if available. Several objects with repeated bright outbursts can be easily spotted in the public lightcurves of the RXTE PCA Bulge Scan (Swank & Markwardt, 2001) [^3] https://asd.gsfc.nasa.gov/Craig.Markwardt//galscan/main.html, the RXTE All-Sky Monitor (Levine et al., 1996) [^4] http://xte.mit.edu; comments at http://xte.mit.edu/XTE/xte_anno.html and/or in MAXI GSC observations (Matsuoka et al., 2009; Suguzaki et al., 2011). We have added some obvious outbursts visible in these surveys to the list (e.g. outbursts from XTE J1709-267 in 2005, 2009, 2018/19, and 2022). The line for defining an ‘outburst’ can be unclear; for simplicity we require an increase of a factor of 10 from the quiescent flux (if known), and $L_{X}>10^{34}$ erg/s.
@@
We separate the objects with many outbursts into separate tables, placed in Appendix B. For the most frequently repeating objects with bright outbursts (of order once per year or more), 4U 1608-52, MXB 1730-33 (the Rapid Burster in Liller 1), Aquila X-1, and GRS 1747-312 (the eclipser in Terzan 6), we use the three public lightcurves above, as well as Swift/XRT public lightcurves (Evans et al., 2007) [^5] https://www.swift.ac.uk/user_objects/, and the INTEGRAL JEM-X Galactic Bulge Monitoring lightcurves (Kuulkers et al., 2007) [^6] http://integral.esac.esa.int/BULGE/ to develop the best possible list of outbursts from 1996 to 2023, along with literature identification of outbursts before 1996. For the objects with the most frequent outbursts, we also record the MJD date of the outburst peak (as best we can tell). We place the outbursts of NGC 6440 X-2 in a separate table in Appendix B as well, as their high frequency requires MJD dates to differentiate them.
@@
We find additional outbursts, not mentioned in the literature, for 12 systems; XTE J1709-267 (4), XTE J1701-407 (1), XTE J1739-285 (2), 1A 1744-361 (2), XMMU J174716.1-281048 (1), AX J1754.2-2754 (3), SAX J1810.8-2609 (1), 4U 1608-52 (3), MXB 1730-335 (aka the Rapid Burster, 52), GRS 1747-312 (37), NGC 6440 X-2 (2), and Aquila X-1 (1), for a total of 109 newly identified outbursts.
@@
$\bullet$ We catalog 37 previously unidentified outbursts from GRS 1747-312 (in the globular cluster Terzan 6). The outburst interval averages 138 days, with significant scatter, but regular enough to indicate that most outbursts are indeed due to a single source. A second transient bursting source, Terzan 6 X-2, has recently been identified in Terzan 6 Saji et al. (2016); van den Berg et al. (2024), but appears not to reach such high fluxes as GRS 1747-312.
In §3 we use the information in these tables to draw some general conclusions about our sensitivity to X-ray outbursts, and apply this information to consideration of interesting X-ray sources near the Galactic Center.
@@
IGR J00291+5934 2008/8 2008/10 6.1e-10 - 1.7 1.3e36 4.2 Lewis et al. (2010)
IGR J00291+5934 2015/8 2015/8 5.6e-10 1e22 1.7 1.2e36 4.2 De Falco et al. (2017)
@@
GS 0836-429 2003/1 2003/5 1.0e-9 2.7e22 1.5 1.0e37 $<$9.2 Aranzana et al. (2016)
GS 0836-429 2003/9 2004/6 1.3e-9 2.7e22 1.5 1.3e37 $<$9.2 Aranzana et al. (2016)
MAXI J0911-655 2016/2 2023/10 1.7e-10 2.7e21 1.27 1.9e36 9.5 Sanna et al. (2017b); Ng et al. (2021a),MAXI
XTE J0929-314 2002/4 2002/4 6.7e-10 7.6e20 1.7-2 $>$2.9e36 $>$6 Galloway et al. (2002); Juett et al. (2003)
MAXI J1421-613 2014/01 2014/1 1.9e-9 4.8e22 2.1 2e36 3 Serino et al. (2015); Nobukawa et al. (2020)
SRGA J144459.2-604207 2024/2 2024/3 2.7e-9 2.9e22 1.8 2.3e37 8.5 Molkov et al. (2024); Ng et al. (2024)
Cen X-4 1979/05 1979/5 9.6e-8 9e20 - 1.9e37 1.3 Kaluzienski et al. (1980); Kuulkers et al. (2009)
Table 1: Published outbursts of verified neutron star LMXBs. Max $F_{X}$ is highest reported flux in each outburst, transformed to 2-10 keV, generally using the power-law spectral index and $N_{H}$ noted. We assume 1 Crab is $F_{X}$(2-10)=$2\times 10^{-8}$ erg s-1 cm-2. Peak observed intrinsic $L_{X}$ in 2-10 keV, for the assumed distance (listed). The full Table is given in the appendix (for arXiv), or online (for journal); a portion is reproduced here.

### v2

@@ Abstract
Many X-ray binaries are transiently accreting. Having statistics on their recurrence times is helpful to address questions related to binary evolution and populations, as well as the physics of binary systems. We compile a catalog of known outbursts of 87 transient neutron star (identified through bursts or pulsations) low-mass X-ray binaries, until mid-2025. Most outbursts are taken from the literature, but we also identify some outbursts from public X-ray monitoring lightcurves. We find 109 outbursts not previously identified in the literature; most are from the frequent transients GRS 1747-312, and the Rapid Burster MXB 1730-335, though we suspect that two outbursts from Liller 1 may be from another transient, besides the Rapid Burster. We also find new outbursts for 10 other systems, and verify substantial quiescent intervals for XMM J174457-2850.3, XMMU J174716.1-281048, and AX J1754.2-2754. Outburst detection has been relatively efficient since 1996 for outbursts above $F_{X}$(2-10 keV)$=3\times 10^{-10}$ ergs cm-2 s-1. While several systems have many known outbursts, 40 of the 87 systems we track have zero or one recorded outburst between 1996 and 2023. This suggests that many faint Galactic Center X-ray binaries may be neutron star X-ray binaries, though we cannot completely rule out the proposition that most neutron star X-ray binaries undergo frequent outbursts below all-sky monitor detection limits.
@@ 1 Introduction
The total number of mass-transferring neutron star low-mass X-ray binaries in the Galaxy is thought to be of the order of $10^{4}$ (Portegies Zwart et al., 1997; Kiel & Hurley, 2006; Jonker et al., 2011; van Haaften et al., 2015). However, only a small fraction of these systems have been detected, typically during outbursts in which their X-ray luminosity increases by a factor of $10^{3}$$10^{7}$ compared to quiescence. Understanding the full population of these systems, and their binary evolution, is helpful for us to understand the physics of binary systems (e.g. Podsiadlowski et al., 2002; Postnov & Yungelson, 2014; Van et al., 2019), and also to understand the products of their evolution, such as millisecond pulsars (Alpar & Shaham, 1985; Verbunt, 1993; Tauris & Savonije, 1999; Tauris et al., 2012), which may produce the excess gamma-rays seen from the Galactic bulge (Wang et al., 2005; Gonthier et al., 2018). One question of current interest is the nature of X-ray sources of intermediate luminosity ($L_{X}$(2-10 keV)$\sim 10^{31}-10^{33}$ erg/s) detected near the Galactic Centre (Muno et al., 2005; Hailey et al., 2018; Mori et al., 2021; Maccarone et al., 2022). The objects in this region that show no outbursts have been argued to be mostly black hole LMXBs, based on an argument that neutron star LMXBs do not have long enough recurrence times. Having detailed statistics on X-ray outbursts from neutron stars can be helpful in beginning to answer questions such as these.
@@
A number of catalogs of X-ray binaries have been produced over the years. The catalogs of Liu et al. (2001, 2007) have been particularly well-cited, but are now 17 years old. Ritter & Kolb (2003) have an extensive and updated catalog of X-ray binaries and cataclysmic variables with known periods, but do not include objects without known (or suggested) orbital periods. More recent catalogs have been published of black hole X-ray binaries, one focused on compiling detailed lightcurves (Tetarenko et al., 2016b), and one focusing on multiwavelength studies (Corral-Santana et al., 2016). We are interested in the particular problem of understanding the frequency of X-ray outbursts from neutron star LMXBs, but found that we lacked a catalog to answer this question. The catalogs of Avakyan et al. 2023 and Fortin et al. (2024) were released when this work was largely complete, and while they are very useful, they also do not list outbursts. Yan & Yu (2015) and Lin et al. (2019) construct tables of outburst frequency for a number of X-ray binaries with known orbital periods, but these works only include 14 and 19 NS LMXBs respectively. Yan & Yu (2015) do include a detailed listing of the outbursts they include, which is helpful. We identify several more outbursts from 4U 1608-52, XTE J1709-267, MXB 1730-33, XTE J1739-285, 1A 1744–361, SAX J1748.9-2021, and Aquila X-1 in the time range these authors cover, typically fainter outbursts.
@@ 2 Catalog
We omit 41 NS LMXBs in the MINBAR catalog that appear to be persistent. Identifying whether a NS LMXB is persistent can be tricky. Seven known NS transients have had outbursts lasting over a decade (EXO 0748-676, KS 1731-260, XB 1733-30, IGR J17062-6143, HETE J1900.1-2455, 1H 1905+000, and 4U 2129+47), and it is possible that systems we list as persistent will turn off in the future. Other systems are persistently active at a low level, then brighten substantially for weeks or months (e.g. 4U 1705-32, Negoro et al. 2021a; in’t Zand et al. 2007). UW CrB (1E 1603.6+2600) at a distance of $4.6^{+1.1}_{-2.5}$ kpc (Gaia parallax, Bailer-Jones et al., 2021) appears unusually faint for an actively accreting system, at $L_{X}\sim 3\times 10^{33}$ erg s-1, but this is thought to be due to extreme obscuration of its intrinsic X-ray luminosity (see e.g. Hakala et al., 2005). We list the systems we consider persistent, verified bursting or pulsing NS XRBs in Table 10, in Appendix D.
@@
We omit a few transient bursters; the position of XB 1940-04 (Murakami et al., 1983) is very poorly known. The eccentric X-ray binary Cir X-1 (or 3A 1516-569) has outbursts that typically reflect its orbital period, not disk outbursts (Murdin et al., 1980). During the period from MJD=55000–58500, it was unusually under-luminous, and showed occasional outbursts that were not at exact multiples of its orbital period, suggesting outbursts from a disk. We catalog those outbursts in Appendix B, but due to Cir X-1’s differences in accretion mode compared to other LMXBs, we omit it from most of our compiled analyses. We exclude 1A 1742-289 (or A1742-289), as it is not clear that the burst activity seen from MXB 1743-29 in the Galactic Center in 1976 (Lewin et al., 1976b) can be attributed to this radio-bright (Davies et al., 1976) source (we note that this is definitely not the same object as AX J1745.6-2901; Kennea & Skinner 1996). This leaves 70 sources from the MINBAR catalog (71 when including Cir X-1).
We add to the MINBAR list the known recent bursters111as updated by Jean in ’t Zand; https://burst.sci.monash.edu/wiki/index.php?n=MINBAR.SourceTable 4U 1730-22, MAXI J1816-195, MAXI J0556-332, MAXI J0911-655, Terzan 6 X-2 (see Appendix § B), and SRGA J144459.2-604207  [^2] We include SRGA J144459.2-604207 in our list for completeness, but our statistical calculations include data only through 2023..
@@
We also add known transient accreting millisecond pulsars that are not known bursters, from the Patruno & Watts (2021) catalog and more recent literature, XTE J0929-314, NGC 6440 X-2, XTE J1751-305, Swift J1756.9-2508, XTE J1807-294, IGR J17494-3030, IGR J16597-3704, and MAXI J1957+032. Finally, we include the transient slow (2.14 Hz) accreting pulsar GRO J1744-28. (We also list the two persistent LMXBs that are pulsars in Table 10.) This gives us a total of 86 transient neutron star low-mass X-ray binaries, 87 if including Cir X-1.
@@
We performed literature searches for this list of transients, and identify the known X-ray outbursts from each system. For each outburst, we attempt to determine the peak observed 2-10 keV X-ray luminosity. This will have some significant uncertainty, since not every outburst is caught at its peak, some outbursts are observed only in other energy ranges, and many distances are not well-known. We attempt to use the best-fit spectrum to convert measurements to 2-10 keV fluxes, but we only attempt this with simple power-law spectra (we do not try to convert complex Comptonized models). Where information is lacking, we use the best literature $N_{H}$ estimate and an assumed photon index of 2.1 to make conversions, or use literature estimates of the 2-10 keV flux in mCrab (assuming 1 Crab = $2\times 10^{-8}$ ergs cm-2 s-1 ). A few bursters have recorded bursts without any recorded outbursts (Gotthelf & Kulkarni, 1997; Cornelisse et al., 2002b); we list these in Table 4 in Appendix C.
@@
We present the catalog of known outbursts of neutron star LMXBs in Table 1, for which we provide a sample here. The full table is in (Table A) in Appendix A). We include a line for each recorded outburst, with a literature reference for each outburst if available. Several objects with repeated bright outbursts can be easily spotted in the public lightcurves of the RXTE PCA Bulge Scan (Swank & Markwardt, 2001) [^3] https://asd.gsfc.nasa.gov/Craig.Markwardt//galscan/main.html, the RXTE All-Sky Monitor (Levine et al., 1996a) [^4] https://heasarc.gsfc.nasa.gov/docs/xte/asm_products.html and/or in MAXI GSC observations (Matsuoka et al., 2009; Suguzaki et al., 2011) [^5] http://maxi.riken.jp/top/index.html. We have added some obvious outbursts visible in these surveys to the list (e.g. outbursts from XTE J1709-267 in 2005, 2009, 2018/19, and 2022). The line for defining an ‘outburst’ can be unclear; for simplicity we require an increase of a factor of 10 from the quiescent flux (if known), and $L_{X}>10^{34}$ erg/s.
@@
We separate the objects with many outbursts into separate tables, placed in Appendix B. For the most frequently repeating objects with bright outbursts (of order once per year or more), 4U 1608-52, MXB 1730-33 (the Rapid Burster in Liller 1), Aquila X-1, GRS 1747-312 (the eclipser in Terzan 6), and Cir X-1, we use the three public lightcurves above, as well as Swift/XRT public lightcurves (Evans et al., 2007) [^6] https://www.swift.ac.uk/user_objects/, and the INTEGRAL JEM-X Galactic Bulge Monitoring lightcurves (Kuulkers et al., 2007a) [^7] http://integral.esac.esa.int/BULGE/ to develop the best possible list of outbursts from 1996 to 2023, along with literature identification of outbursts before 1996. For the objects with the most frequent outbursts, we also record the MJD date of the outburst peak (as best we can tell). We place the outbursts of NGC 6440 X-2 in a separate table in Appendix B as well, as their high frequency requires MJD dates to differentiate them.
@@
We find additional outbursts, not mentioned in the literature, for 12 systems; XTE J1709-267 (4 outbursts), XTE J1701-407 (1), XTE J1739-285 (2), 1A 1744-361 (2), XMMU J174716.1-281048 (1), AX J1754.2-2754 (3), SAX J1810.8-2609 (1), 4U 1608-52 (3), MXB 1730-335 (aka the Rapid Burster, 52), GRS 1747-312 (37), NGC 6440 X-2 (2), and Aquila X-1 (1), for a total of 109 newly identified outbursts. We also identify 42 additional outbursts from Cir X-1 not mentioned in the literature between 2006 and 2018, though we list these separately as their nature is somewhat ambiguous.
@@
$\bullet$ We catalog 37 previously unidentified outbursts from GRS 1747-312 (in the globular cluster Terzan 6). The outburst interval averages 138 days, with significant scatter, but regular enough to indicate that most outbursts are indeed due to a single source. A second transient bursting source, Terzan 6 X-2, has recently been identified in Terzan 6 (Saji et al., 2016; van den Berg et al., 2024), but appears not to reach such high fluxes as GRS 1747-312.
In §3 we summarize some useful information on monitoring instruments, use the information in these tables to draw some general conclusions about our sensitivity to X-ray outbursts, and apply this information to consideration of interesting X-ray sources near the Galactic Center.
@@
IGR J00291+5934 2008/8 2008/10 6.1e-10 - 1.7 1.3e36 4.2 Chakrabarty et al. (2008); Lewis et al. (2010)
IGR J00291+5934 2015/8 2015/8 5.6e-10 1e22 1.7 1.2e36 4.2 Cummings et al. (2015a); De Falco et al. (2017)
@@
Table 1: Published outbursts of verified neutron star LMXBs. Max $F_{X}$ is highest reported flux in each outburst, transformed to 2-10 keV, generally using the power-law spectral index and $N_{H}$ noted. We assume 1 Crab is $F_{X}$(2-10 keV)=$2\times 10^{-8}$ erg cm-2 s-1. Peak observed intrinsic $L_{X}$ in 2-10 keV, for the assumed distance (listed). Reference to an instrument (e.g. MAXI) indicates we have made use of that instrument’s archive to generate some information for that line. The full Table is Table A in Appendix A; a portion is reproduced here.
@@ 3 Discussion
3.1 Monitoring instruments
We first review key information about some relevant monitoring instruments.
The first effective all-sky X-ray monitor was a scintillation 3-12 keV X-ray detector on Vela 5B, which flew from 1969-1979. Its limited collimation and thus high background limited its sensitivity to of order 0.3 Crabs (Priedhorsky & Holt, 1987). Ariel V (1974-1980) had a slightly more sensitive (perhaps 0.2 Crabs) 3-6 keV all-sky monitor, operated as a pinhole camera (Holt, 1976). In 1987-1991, Ginga’s ASM operated from 1-20 keV with sensitivity down to 50 mCrab (Tsunemi et al., 1989).
RXTE/ASM monitoring produced a dramatic increase in detected outbursts, from January 1996 until 2009-2011 (Levine et al., 1996a) [^8] The RXTE/ASM sensitivity decayed from 2009 until RXTE was turned off in 2011.. RXTE/ASM achieved roughly 15 mCrab ($F_{X}=3\times 10^{-10}$ ergs cm-2 s-1, 3-sigma) sensitivity per day, in 2-10 keV, for known sources. The RXTE/ASM team could identify new sources as faint as 50 mCrab with one day’s data, while integrating for a week could delve down to 7 mCrab, in uncrowded regions (Bradt et al., 2001). Indeed, detections from 1996 on did reach $3\times 10^{-10}$, though the completeness limit was higher, perhaps $10^{-9}$ ergs cm-2 s-1, or 50 mCrab. The RXTE/ASM covered most sources daily, except for about 30 days each year when the Sun was nearest for sources in the ecliptic.
From 1996 to 2002, in each spring and fall, BeppoSAX conducted regular monitoring of the Galactic Bulge with its Wide-Field Camera (WFC), with a 40o by 40o field of view, 5’ angular resolution, and sensitivity of about 15 mCrab in 2-28 keV (Heise et al., 1999; in’t Zand et al., 2004). Monitoring of the Galactic Center was performed weekly for two months in the spring and fall (Heise et al., 1999). Jager et al. (1997) estimated the typical observation sensitivity as 10 mCrab at high galactic latitudes, and 30 mCrab at the Galactic Center. BeppoSAX’s WFC discovered a large number of faint ($L_{X}<10^{37}$ erg/s) X-ray transients in the Galactic Bulge, and also several X-ray bursts from transients that were too faint for the WFC to detect individually (Cornelisse et al., 2002b).
From 1999 to 2011, RXTE also conducted monitoring of the Galactic Bulge using slews of the PCA instrument (Swank & Markwardt, 2001). These Bulge scans could detect new uncrowded sources down to 0.5 mCrab (Swank & Markwardt, 2001), but in practice, in this crowded region, they reached a sensitivity of, at best, 5 mCrab, or $10^{-10}$ ergs cm-2 s-1 in 2.5-10 keV. Notably, the RXTE/PCA was able to detect several faint ($\sim$5 mCrab) outbursts of NGC 6440 X-2 in 2009-2011 (Heinke et al., 2010), but these were only barely visible in the Bulge Scan lightcurves. The PCA Bulge scans initially covered about 250 square degrees twice weekly (Markwardt et al., 2002b), extended to 500 square degrees in 2003 (Markwardt, 2006), and for 2009-2011 added the Aquila region, to cover over 700 square degrees (Markwardt & Swank, 2010). These scans were paused for the 2 months when the Sun is closest to the Galactic Center (Swank & Markwardt, 2001).
The INTEGRAL satellite (launched 2002, decommissioned in Feb. 2025) has two relevant coded-aperture instruments: IBIS (Ubertini et al., 2003) has peak sensitivity in the 20-60 keV range with a 20 degree field of view (at half sensitivity) and 12’ angular resolution, while JEM-X (Lund et al., 2003) observes in 3-35 keV, with a 7.5 degree field of view and 3.4’ angular resolution. The INTEGRAL Galactic Bulge hard X-ray monitoring program has been particularly productive, with 5-15 mCrab sensitivity per 1800 s observation (every three days), and summed detections reaching down to 2.4 mCrab (Kuulkers et al., 2007a). INTEGRAL could observe the Bulge only 1/3 of the year, for 2 months in each spring and fall. Thirteen of our transients have principal names starting with IGR (only XTE matches this number), suggesting INTEGRAL’s impact. There appears (Fig. 2) to be a small decline in detections in the past few years, which may be due to declining sensitivity of INTEGRAL’s IBIS/ISGRI instrument. [^9] The Integral Galactic Bulge Monitoring Program’s website, http://integral.esac.esa.int/BULGE/, presents 18-40 keV IBIS/ISGRI lightcurves for sources, which show increasing noise from 2016 and stop in 2018.
Swift/BAT’s hard X-ray (15-50 keV) monitoring program has operated since 2006, with a daily 16 mCrab detection limit (Krimm et al., 2013). Swift’s XRT instrument (narrow-field imaging in 0.3-10 keV) has also been vital in tracking transients, due to its flexible scheduling, ease of ToO requests, good sensitivity, and simplicity of data analysis (e.g. Evans et al., 2007).
The MAXI instrument (on the International Space Station) has two camera groups, of which the Gas Slit Cameras, or MAXI/GSC (Mihara et al., 2011; Sugizaki et al., 2011) has been significantly more productive. MAXI/GSC has a field of view of 1.5o by 160o, and monitors $\sim$85of the sky every 92 minutes, reaching 95% of the sky each day (Negoro et al., 2016b). MAXI/GSC has monitored the sky since 2009 daily in 2-20 keV, with the best sensitivity in 3-10 keV. Typical sensitivities for 4-day integrations reach 7 mCrab (Negoro et al., 2016b). Positions of isolated sources are typically determined to 0.1o, but MAXI/GSC’s power to resolve the Galactic Center is limited. MAXI’s efficient and automated nova search algorithms (Negoro et al., 2016b) have been highly productive, identifying transients which can then be followed up by other missions to get precise positions.
The SRG spacecraft scans the sky twice per year, and contains two instruments: ART-XC in 4-30 keV, and eROSITA in 0.2-8 keV (Sunyaev et al., 2021). eROSITA completed 4 of 8 planned all-sky surveys (2019-2022), with sensitivity per survey (200 s exposure) reaching $F_{X}=10^{-13}$ ergs cm-2 s-1 in 2-4 keV (Sunyaev et al., 2021), before the Ukraine war terminated Russian/German cooperation. ART-XC continues operating, and reaches 1 mCrab sensitivity in 4-12 keV each survey sweep (Sunyaev et al., 2021). ART-XC detected a new X-ray binary, SRGA J144459.2-604207 (Molkov et al., 2024), during the preparation of this catalog.
Two new instruments capable of transient LMXB identifications began surveying the sky within the last year. The Einstein Probe is a Chinese soft X-ray mission launched on January 9, 2024. Its Wide-Field X-ray Telescope (WXT) monitors the entire night sky on a daily basis, using lobster-eye optics to achieve a 3600 square degree instantaneous field of view, a spatial resolution of 5’, and a nominal 0.5-4 keV sensitivity of 0.8 mCrab in 1000 s (Yuan et al., 2022). The Follow-up X-ray Telescope (FXT) has 30” angular resolution, and sensitivity similar to Swift/XRT, reaching $F_{X}$(0.5-2 keV)$=10^{-14}$ erg/s in 1500 s (Yuan et al., 2022).
Finally, the Space Variable Objects Monitor (SVOM) is a joint Chinese/French mission launched on 22 June 2024, with a science focus on gamma-ray bursts. Its wide-field hard X-ray (4-150 keV) detector, ECLAIRS, has a sensitivity of 50 mCrab per orbit, with a PSF of 52’ (Godet et al., 2014). It also hosts a narrow-field soft X-ray (0.2-10 keV) detector, MXT, with angular resolution of 4.5’ and a sensitivity reaching to $F_{X}$(0.2-10 keV)$10^{-12}$ with 10 ks of integration. SVOM’s mission plan of generally pointing away from the Galactic Plane will tilt its science toward GRBs, but it is likely to also detect LMXB outbursts (Wei et al., 2016).
Both INTEGRAL/IBIS and Swift/BAT are sensitive to harder X-rays than the 2-10 keV range we focus on here. Thus, they will be more likely to detect LMXBs in hard states, vs. soft states at the same 2-10 keV flux. The ratio of Crab-normalized Swift/BAT 15-50 keV vs. MAXI/GSC 2-10 (or 4-10) keV transient NS LMXB fluxes is about 1 in the hard state, and often 0.1 in the soft state (e.g. Asai et al. 2015, Bahramian et al. 2014; cf. Tetarenko et al. 2016b for BH LMXBs, which can reach lower ratios in the soft state). Most transient NS LMXBs observed near the limit of all-sky monitor detection ($\sim$5 mCrab) at distances of 10 kpc (thus, $L_{X}$(2-10)$\sim 10^{36}$ erg/s) are found to be in the hard spectral state (Maccarone & Coppi, 2003; Asai et al., 2015). However, INTEGRAL/IBIS and Swift/BAT would have a 2-10 keV $L_{X}$ detection threshold roughly 10 times higher, for soft-state transients.